OpenStudy (anonymous):

a bag consists of 4 red, 5 white, and 3 blue marbles. find the probability of obtaining a white or a blue marble in a single draw.

OpenStudy (anonymous):

total=4+5+3=12 marbles P(white or blue) = (total white or blue marbles) / total P(w or b) = (5+3) / 12 P(w or b) = 8/12 P(w or b) = 2/3 2/3, or 67%.

OpenStudy (anonymous):

P(white)=5/12 P(blue)=3/12 P(total)=(5+3)/12=8/12=2/3